OpenStudy (anonymous):

How would I solve 5x - x = x + 3x?

OpenStudy (anonymous):

combine like terms...

OpenStudy (anonymous):

5x-x = x + 3x 4x= x+3x 4x = 4x

OpenStudy (radar):

I don't think it is a single value solution. Other words, x can take on an infinite number of values.

OpenStudy (anonymous):

yes because equal sign still would hold.